A sample of 50 socks is taken. What is the probability that the mean length of the sample is shorter than 18 cm?


1
Expert's answer
2021-05-24T03:00:01-0400

"n=50,



\mu=20"


"SE=\dfrac{\sigma}{\sqrt{n}}=\dfrac{10.5}{\sqrt{50}}=\dfrac{10.5}{7.07}=1.48"

 Probability that the mean length of the sample is shorter than 18 cm is-


"P(X<18)=P(z<\dfrac{18-20}{1.48}"



"P(X<18) = P(Z < -1.35)= 0.0885"


THE PROBABILITY IS 0.0885